# Constraint Learning for Parametric Point Cloud

### v3

@@ Abstract
Parametric point clouds are sampled from CAD shapes, and have become increasingly prevalent in industrial manufacturing. However, most existing point cloud learning methods focus on the geometric features, such as developing efficient convolution operations, overlooking the important attribute of constraints inherent in CAD shapes, which limits these methods’ ability to comprehend CAD shapes fully. To address this issue, we analyzed the effect of constraints, and proposed its deep learning-friendly representation, after that, the Constraint Feature Learning Network (CstNet) was developed to extract and leverage constraints. Our CstNet includes two stages. Stage 1 extracts constraints from B-Rep data or point cloud. Stage 2 leverages coordinates and constraints to enhance the comprehension of CAD shapes. Additionally, we built up the Parametric 20,000 Multi-modal Dataset for the scarcity of labeled B-Rep datasets. Experiments demonstrate that our CstNet achieved state-of-the-art performance on both public and proposed CAD shape datasets. To the best of our knowledge, CstNet is the first constraint-based learning method tailored for CAD shape analysis.
Project page: https://cstnetwork.github.io/
@@ 1 Introduction
Parametric point clouds are derived from parametric templates, which consist of primitives, constraints, and dimensional parameters. By specifying all parameters, the templates can be instantiated into CAD shapes, as illustrated in Fig. 1. Parametric point clouds provide detailed information that is essential for various engineering applications, such as model design and product machining.
In recent years, deep learning methods have emerged as powerful tools for point cloud analysis, with models such as PointNet [33], PointNet++ [34], and PointCNN [24] demonstrating significant success across various tasks. These models are primarily designed and evaluated on graphic objects, such as animals, trees, and other free-form shapes, where shape comprehension relies on geometric features visible in contours. However, some CAD shapes, despite exhibiting minimal visual differences, serve distinct purposes due to variations in their functional regions. This similarity in appearance poses challenges in distinguishing them based on geometric features. Notably, variations in the functional regions often introduce differing constraints, as these regions need to fit closely with other parts to perform functions such as transmission and positioning. Thus, distinguishing them from a constraint-based perspective becomes more feasible, as illustrated in Fig. 1.
@@
Figure 1: Motivation. Some similar CAD shapes serve different functions. The eccentric wheel has a rotation axis that is offset from its outer cylindrical surface, facilitating precise positioning or tension adjustment. In contrast, the flywheel has a coaxial rotation axis with its outer cylindrical surface and is used to store rotational kinetic energy. Although these shapes are challenging to distinguish visually, they can be easily differentiated from the perspective of constraints.
Additionally, the limited availability of B-Rep datasets poses barriers to deep learning on CAD shapes. Most existing CAD shape datasets consist of mesh files, such as MCB [16] and ESB [12]. While mesh files could approximate the appearance of CAD shapes, lack crucial boundary information. In contrast, B-Rep data [19, 57] serves as the native representation of CAD shapes and is therefore more suitable for dataset construction. However, labeled B-Rep datasets remain relatively scarce, for example, FabWave [1] includes only 2,133 B-Rep files. Although the ABC [18] contains a large number of B-Rep files, it remains unlabeled.
To facilitate the constraint feature learning of CAD shapes, we conducted studies on deep learning methods and dataset development. Traditional constraint definitions are not well-suited for deep learning, as they vary across constraint types and involve multiple possible combinations to represent primitives’ relations. To address this challenge, we proposed a novel constraint representation as point-wise local attributes. Afterward, we developed the CstNet for constraint extraction and feature learning, which consist of two stages. Stage 1 is built for constraint acquisition, if B-Rep data is available, the openCASCADE Technology (OCCT) is used to extract constraints, otherwise, the constraint prediction from point cloud (CST-PCD) module is adopted. The CST-PCD is designed to rely on local features only, enabling it to generalize to unseen datasets after pre-training. Stage 2 leverages both point cloud and constraints for constraint feature learning, which employs both point-level and feature-level attention for effective feature extraction, enhancing the comprehension of CAD shapes. Finally, we built up a multi-modal classification dataset Parametric 20000, which contains B-Rep, Mesh, and point cloud. Given the current scarcity of labeled CAD shape datasets, Parametric 20000 offers researchers an expanded resource for their work. Our contributions are threefold:
  • We introduced the deep learning-friendly expression of constraints.
  • We designed the CstNet for constraint extraction and constraint feature learning.
  • We collected and curated the Parametric 20000 dataset, which consists of mesh, point cloud, and B-Rep data.
Our CstNet has been validated across various tasks, on the mechanical classification dataset MCB [16], our model achieved an overall accuracy improvement of 2.98 % compared to the state-of-the-art, and for mechanical segmentation dataset 360 Gallery [19], that is 6.74 %.
@@ 2 Related Work
For the published deep learning methods focusing on CAD shapes are very few, and even fewer aimed at parametric point clouds, this section will provide an overview of existing point cloud learning methods as well as CAD shapes-oriented models. By integrating insights from the above areas, a comprehensive review of the relevant works could be presented.
2.1 Deep learning on point clouds
Point cloud deep learning can be categorized into three types based on data processing approaches: view-based, voxel-based, and point-based methods.
View-based methods typically project the point cloud from multiple views, the obtained figures are then fed in convolutional neural networks (CNNs) for feature extraction [38, 15, 6, 55, 54, 43]. However, these methods face significant challenges when applied to large-scale scanned data, primarily due to the difficulty of covering entire scenes from single-point perspectives, and the choice of projection views significantly impacts recognition performance. Besides that, it cannot be extended to 3D data segmentation or reconstruction easily.
Voxelization-based methods convert point clouds into 3D voxels, and then employ 3D CNNs for feature extraction [30, 48, 36, 41, 17, 20]. However, the size and orientation of the voxels can affect computational efficiency and memory usage. The voxel grid quantization process may lead to geometric detail loss, and the inherent sparsity of point clouds may result in many useless voxels.
Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [33, 34, 24, 10, 53, 2, 39, 27, 59, 60, 9, 7, 51, 50, 49, 35, 28, 26, 42, 47, 4, 58]. Typically, symmetric functions and k-nearest neighbors (KNN) are employed for feature extraction. However, the unordered nature and transformation invariance of point clouds pose challenges for efficient feature extraction.
2.2 Deep learning on CAD shapes
CAD shapes deep learning can be categorized into three types based on the input data format: B-Rep-based learning, mesh-based learning, and sketch-based learning.
@@ 2.2 Deep learning on CAD shapes
B-Rep-based methods typically involve B-Rep data conversion and feature extraction. Since B-Rep data cannot be directly fed into deep learning models, the challenge lies in converting B-Rep into deep learning compatible format, with graph representation commonly used. These approaches are commonly applied to classification [13], segmentation [19, 5], assembly [14, 44], engineering sketch generation [11], and CAD operation sequence prediction [46, 61].
Mesh-based learning methods use triangle models as input, such as STL and OBJ. Compared to B-Rep, mesh representations lack boundary details, which prevents these methods from leveraging data conversion for improved performance. As a result, contributions often rely on mechanical datasets provided by the authors. These methods are primarily used for classification [29], retrieval [3], segmentation [40], and B-Rep generation [8].
Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [8, 45, 31, 37, 25], 3D models [21, 22, 56], or CAD operation sequences [52]. Most of these models are generative, and frequently utilize transformer modules. A key challenge is incorporating the regularity and connectivity of primitives, which are typically overlooked in graphic object analysis.
@@
While there has been substantial progress in above domains, the engineering applications of parametric point cloud analysis demand further innovation to ensure accuracy and robustness.
@@ 3 Method
3.1 Problem Statement
Constraints are crucial for CAD shapes, can they be utilized to enhance deep learning methods’ performance? To answer above question, we designed the following validation experiments. The objective is to compare the prisms shown in Fig. 2 with cuboids. In each experiment, only prisms with a specific angle are selected, creating a binary classification task between prisms and cuboids. Consequently, eight independent experiments are conducted to cover all angles. As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them. The constraint-aware model is built on PointNet++ backbone, this model predicts constraints and then adopts it for classification. The constraint representation is defined as the relation between the point attached primitive and a reference plane, more details in supplementary.
@@ 3.1 Problem Statement
The experimental results are presented in Fig. 3. For prisms with angle of 50°50°50  ° and 89°89°89  °, the constraints have little improvement on the results. The reason lies in the 50°50°50  ° prism differs substantially from the cuboid, while the 89°89°89  ° prism is too similar to the cuboid. For other circumstances, incorporating constraints improves classification accuracy significantly, demonstrating that the constraint enhanced the model’s comprehension of CAD shapes.
@@ 3.2 Constraint Representation
Constraint should be defined as point-wise attribute for point cloud, analogous to normal vector, as point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation.
Following the constraint definitions in CAD software is not applicable. However, once the constraint is applied, the positional relations between primitives become fixed. Therefore, if the type of primitives and their relations are determined, the constraints on CAD shapes can be defined comprehensively.
The point-wise primitive type can be represented as the type of the primitive to which the point is attached, using one-hot encoding.
The positional relation is defined as the combination of main axis direction and adjacency. Considering that the direction of primitives’ main axis is unique, as shown in Fig. 4, relations such as parallel and vertical could be derived from the direction of their main axes. However, some dimensional relations can not, such as the cylinders’ radius equivalence. Thus, additional data is required to supplement this information.
Figure 4: Main axis of primitives. Plane: normal vector. Cylinder, cone, and sphere: rotation axis.
Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitations on the distance and size of primitives. For example, in a cuboid, the adjacency and normal of the six faces ensure that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one. Therefore, if the adjacency is combined with main axis directions, the positional relation can be determined comprehensively.
Adjacency represents the connection between primitives and cannot be expressed as point-wise attribute directly. The edges are key indicators of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges. For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by two-channel one-hot vector.
In summary, each composition of constraints in parametric point cloud is illustrated in Fig. 5, and each point with constraint is expressed as (Coordinate, Primitive Type, Main Axis Direction, Adjacency).
@@
3.3 Method Overview
Our CstNet and module details are depicted in the top and bottom sides of Fig. 6 respectively. It should be noted that the CST-PCD module requires pre-training, and once the pre-training is completed, its parameters are kept frozen. Given the locality of the constraints expression we designed, the pre-trained CST-PCD module is able to generalize to other unseen CAD shape datasets directly. In this work, the dataset for pre-training CST-PCD module comes from ABC [18].
Figure 6: Overview of CstNet. The top side illustrates the overall architecture of CstNet, comprising two stages. Stage 1 is designed for constraint acquisition. When B-Rep data is available, the CST-BRep module is utilized to extract constraints; otherwise, use the CST-PCD module. Stage 2 performs constraint feature learning, facilitating a deeper understanding of CAD shapes. The bottom side presents the details of module design.
@@ 3.4 Constraint Acquisition
When B-Rep data is available, the CST-BRep module can be used to compute constraints.
For the input B-Rep data, tessellation is performed by OCCT to generate a mesh model, followed by Poisson disk sampling [23] via pymeshlab to ensure a uniform point distribution and control the point count, if point cloud generation is not needed, this step could be skipped.
@@
For adjacency computation, we identify valid edges first. In this study, an edge is considered invalid if the two adjacent primitives are smoothly connected along the edge, this work is accomplished by OCCT. Next, we compute the distance from each point to all valid edges. A point is marked as near an edge if the minimum distance is below a threshold (determined by the model surface area, set to $0.08\times Area/4\pi$ in this paper). For the primitive type and main axis direction of a point, we first determine which primitive it attached, OCCT is then used to analyze the primitive, providing both the primitive type and main axis direction.
@@
In the absence of B-Rep data, such as working with mesh or voxel models, it can be converted into point cloud, after which the CST-PCD module can be used to predict constraints. Since each point’s constraints are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary. The constraint computation of a point $\boldsymbol{p}$ could be divided into the following steps:
Step 1. Identify neighbor points around $\boldsymbol{p}$.
Step 2. From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, these points are referred to as valid points.
Step 3. Fit shapes such as cylinders or planes using valid points, and determine the primitive type with the smallest fitting error.
Step 4. Based on the valid points and the primitive type, calculate the main axis direction and assess whether the point is near an edge.
While the above process can be carried out using traditional methods, it is highly complex, which is why a deep learning model is employed to achieve the same.
@@
The design of CST-PCD module is inspired by the aforementioned calculation process. For Step 1, which involves finding neighbor points, the traditional approach is to use the K Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere. However, for constraint computation, it is more effective to search for neighbors along the shape’s surface. This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig. 7. To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary.
Figure 7: Comparison of KNN and SurfaceKNN. Blue point: center, red point: neighbor.
For Step 2, which involves identifying valid points from neighbors, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding. We designed the point attention layer based on vector self-attention, the relation between points is designed as embedding subtraction, as shown in Eq. 1.
@@
$\boldsymbol{f}^{\prime}_{i}=\sum_{f_{j}\in\mathcal{N}_{i}}\rho\Bigl{(}mlp\bigl{(}\mathrm{Q}(\boldsymbol{f}_{i})-\mathrm{K}(\boldsymbol{f}_{j})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{i}).$ (1)
@@
Where $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication. $\mathrm{Q}(\boldsymbol{f})$, $\mathrm{K}(\boldsymbol{f})$, and $\mathrm{V}(\boldsymbol{f})$ are defined in the same way by concatenating $\boldsymbol{f}$ with the positional encoding $\Delta$ and then feeding it into MLP, where $\Delta$ is defined as $\Delta=\boldsymbol{p}_{i}-\boldsymbol{p}_{j}$.
Step 3 and Step 4 involve determining the Primitive Type and Main Axis Direction, both steps can be accomplished by MLPs. We employed a U-Net-like structure to enhance information utilization. In the CST-PCD module, C-MLP refers to concatenating the input tensors and feeding them into MLPs. The final MLP in this module consists of three parallel MLPs with the same input, where the outputs correspond to Primitive Type, Main Axis Direction, and Adjacency.
@@
The structure of CST-PCD module reveals that only local information is utilized, which aligns with the inherently local nature of the constraint representation we designed. For no global information is incorporated and most CAD shapes consist of a limited set of local geometry, such as plane, cylinder, and sphere, the CST-PCD can generalize to unseen datasets if pre-trained on a large dataset. As a result, our CST-PCD enables the effective use of existing unlabeled B-Rep datasets, making it possible to serve meaningful tasks.
3.5 Constraint Feature Learning

### v5

# Constraint-Aware Feature Learning for Parametric Point Cloud

@@ Abstract
Parametric point clouds are sampled from CAD shapes and are becoming increasingly common in industrial manufacturing. Most existing CAD-specific deep learning methods only focus on geometric features, while overlooking constraints which are inherent and important in CAD shapes. This limits their ability to discern CAD shapes with similar appearance but different constraints. To tackle this challenge, we first analyze the constraint importance via a simple validation experiment. Then, we introduce a deep learning-friendly constraints representation with three vectorized components, and design a constraint-aware feature learning network (CstNet), which includes two stages. Stage 1 extracts constraint feature from B-Rep data or point cloud based on shape local information. It enables better generalization ability to unseen dataset after model pre-training. Stage 2 employs attention layers to adaptively adjust the weights of three constraints’ components. It facilitates the effective utilization of constraints. In addition, we built the first multi-modal parametric-purpose dataset, i.e. Param20K, comprising about 20K shape instances of 75 classes. On this dataset, we performed the classification and rotation robustness experiments, and CstNet achieved 3.52% and 26.17% absolute improvements in instance accuracy over the state-of-the-art methods, respectively. To the best of our knowledge, CstNet is the first constraint-aware deep learning method tailored for parametric point cloud analysis in CAD domain. Our project page with source code is available at: https://cstnetwork.github.io/.
@@ 1 Introduction
Figure 1: Motivation. Some CAD shapes are visually similar but serve different functions. The eccentric wheel has a distance constraint, i.e. its rotation axis is offset from outer cylindrical surface. It is used for precise positioning or tension adjustment. In contrast, the flywheel has a coaxial constraint, i.e. its rotation axis is coaxial with outer cylindrical surface. It is used to store rotational kinetic energy. While these shapes are challenging to distinguish visually, they can be effectively differentiated from constraints.
Parametric point clouds are sampled from CAD shape instances or captured from various industrial scenarios by common digital devices. As illustrated in Fig. 1, CAD shape instances are typically instantiated from parametric templates, which comprise primitives, constraints, and shape parameters. By specifying a series of parameter configurations, a template can be instantiated into a family of CAD shape instances [62]. In industrial domain, many CAD shapes are visual similar but have different constraints in their functional regions. This poses challenges for existing point cloud deep learning methods which based on geometric features only. Thus, a constraint-aware feature learning method for parametric point cloud is fundamental for a wide range of industrial applications, such as engineering model design and product manufacturing.
First, most existing studies [58, 41, 26, 65] focus on primitive regularity (e.g. uniform appearance and distinct boundaries). They usually leverage geometric features only, but ignore those constraints between primitives. Although these methods have been effectively applied in point cloud segmentation and primitive reconstruction, they inherently limit their discrimination capability for CAD shapes that are visually similar yet functionally distinct. While some studies discussed constraints in 2D engineering sketches [8, 40], the considered primitives in 2D sketches are supposed available and generally few in number. Thus, these approaches can be hardly extended to 3D parametric point clouds. Second, most CAD-specific methods rely on specialized datasets with additional labels (e.g. primitive, parameter, and normal labels). However, point clouds are typically acquired via LiDAR or RGB-D cameras, and in most cases only contain coordinate information. It thus restricts the practical applicability of these approaches. Third, there are still numerous challenges in constraint-aware point cloud analysis. Constraints represent the spatial relationships between primitives, but the number and types of primitives cannot be directly obtained from point clouds. Predicting primitives from point clouds also poses challenges. The unordered nature of point clouds results in unordered outcomes, leading to a mismatch between the predicted primitives and those in the constraint labels. In addition, processing relations between every pair of primitives would produce an overwhelming amount of data, since some CAD shapes contain many primitives. Assuming that the constraints have been obtained, effectively leveraging them remains a significant challenge.
To address the aforementioned issues, we transform the traditional constraint representation (e.g. constraints in SolidWorks or UG [3]) into three vectorized sub-features: primitive’s Main Axis Direction (MAD), Adjacency (Adj), and Primitive Type (PT). Given the completeness of traditional constraint representation, these three sub-features can effectively capture constraints. Representing constraints with MAD, Adj, PT converts the relations between primitive pairs into relations between primitives and a common reference, thereby significantly reducing the total number of relations. Moreover, the MAD, Adj, PT can be inferred by local point patches, and the local features of CAD shapes are predominantly composed of primitives such as planes and cylinders. Consequently, a model that relies solely on local features can generalize to unseen datasets after pre-trained, this property eliminates the reliance on constraint labels during practical applications. To mitigate the impact of point cloud unordered nature, we encode MAD, Adj, PT as point-wise features. As a result, each point with constraints in point cloud is represented as (X, Y, Z, MAD, Adj, PT). The aforementioned processing transform the constraints into deep learning-friendly form. After that, we designed the constraint prediction from point cloud (CstPnt) module to obtain MAD, Adj, PT. The CstPnt module is based solely on local features, experiments show that it exhibits strong generalization ability on unseen datasets, and its training data can be easily obtained by processing unlabeled B-Rep data. Following constraint acquisition, we separately extract features from MAD, Adj, PT, and apply attention mechanism to weight these features, thereby facilitating their effective utilization.
We quantitatively evaluate CstNet on ABC [20] and Param20K dataset: constraints prediction, classification, and rotation robustness evaluation. For constraints prediction, our CstPnt module (in CstNet Stage 1) surpasses the state-of-the-art (SOTA) by around 5% and 14% instance accuracy in ABC (test set) and Param20K (unseen dataset), respectively. For classification and rotation robustness on Param20K, CstNet outperforms the SOTA by 3.52% and 26.17% instance accuracy, respectively.
In summary, our contributions are threefold:
  • We introduced a deep learning-friendly vectorized representation and designed the first constraint-aware feature learning network CstNet for effective constraint acquisition and leveraging for parametric point cloud.
  • We built the first parametric-purpose Param20K dataset, which consists of about 20K shape instances of 75 classes. Each instance include mesh, point cloud, and B-Rep data. Given the current scarcity of labeled CAD shape datasets, Param20K offers a solid basis for future studies in parametric CAD domains.
  • We conducted extensive experiments on constraint acquisition, classification, robustness evaluation and ablation studies. The performance of CstNet has a significant improvement compared with SOTA methods which usually focus on geometric data only.
@@ 2 Related Work
General point cloud analysis. View-based methods typically project the point cloud from multiple views, the obtained figures are then fed in convolutional neural networks (CNNs) for feature extraction [42, 17, 7, 60, 59, 47]. How to address information loss during the projection and optimizing viewpoints play a critical role. Voxelization-based methods convert point clouds into 3D voxels, and then employ 3D CNNs for feature extraction [33, 52, 39, 45, 19, 22]. The size and orientation of 3D voxels significantly influences the accuracy of these methods, and information loss occurs during the conversion to voxels. Point-based methods directly use points as input, thereby avoiding information loss associated with view-based and voxel-based methods, it is also the most popular point cloud learning method [36, 37, 27, 12, 57, 2, 43, 30, 64, 66, 11, 9, 55, 54, 53, 38, 31, 29, 46, 51, 5, 63]. The unordered nature and rotational invariance of point clouds make it challenging to employ methods similar to traditional CNNs for feature extraction, and kNN is a commonly adopted approach.
Parametric point cloud analysis. The most extensive application of these methods lies in segmentation and primitive fitting [58, 41, 26, 65]. CAD shapes are typically composed of regular primitives, such as plane, cylinder, cone, and the variety of such elements is limited. Therefore, leveraging primitives’ distinct shape features enables these methods to achieve superior segmentation and primitive fitting outcomes compared to general approaches. These methods generally predicting primitives parameters, and subsequently fed into downstream modules, these modules may incorporate specific primitive fitting algorithms. Despite their superior ability to exploit the shape features of individual primitives, these methods overlooked the constraints between primitives. This oversight limited their ability to distinguish CAD shapes that share similar appearances but serve different functions. Moreover, the primitive parameters predicting procedure necessitates the corresponding parameter labels. In practical, when sampling point clouds from CAD shapes, only coordinates can be obtained in most cases, which limits the applicability of these methods.
@@
Other CAD-specific studies. B-Rep-based methods typically involve B-Rep data conversion and feature extraction, the challenge lies in converting B-Rep into deep learning compatible format. These approaches are commonly applied to classification [15], segmentation [21, 6], assembly [16, 48], engineering sketch generation [13], and CAD operation sequence prediction [50, 67]. Mesh-based learning methods use triangle models as input, such as STL and OBJ, contributions often rely on mechanical datasets provided by the authors. These methods are primarily used for classification [32], retrieval [4], segmentation [44], and B-Rep generation [10]. Sketch-based learning takes hand-drawn sketches as input, with outputs typically being engineering sketches [10, 49, 34, 40, 28], 3D models [23, 24, 62], or CAD operation sequences [56]. A key challenge is incorporating the regularity and connectivity of primitives, which is typically overlooked in graphic object analysis.
@@
While there has been substantial progress in CAD-specific studies, further innovation on constraint learning is demanded to ensure higher accuracy and robustness.
@@ 3 Method
3.1 Is Constraint Important?
We aim to utilize constraints to distinguish visually similar but functionally distinct CAD shapes. To answer whether constraints could enhance deep learning methods’ performance, we designed the following validation experiments. The objective is to compare the prisms shown in Fig. 2 with cuboids. In each experiment, only prisms with a specific angle are selected, creating a binary classification task between prisms and cuboids. Consequently, eight independent experiments are conducted to cover all angles. As the prism’s angle approaches 90°90°90  °, its resemblance to cuboids increases, making it progressively more challenging to distinguish them. The constraint-aware model is built on PointNet++ backbone, which predicts constraints and then adopts it for classification. The constraints are represented as the relation between the point attached primitive and a reference plane, more details in supplementary.
@@ 3.1 Is Constraint Important?
From results presented in Fig. 3, for prisms with angles from 50°50°50  ° to 87°87°87  °, incorporating constraints improves classification accuracy significantly, while the 89°89°89  ° prisms are too similar to the cuboids, all methods cannot distinguish them.
@@ 3.2 Constraint Representation
In general, constraints should be defined as point-wise attribute for point clouds, analogous to normal vector. As point cloud is inherently unordered, assigning constraints at the point-wise level ensures that their representation is independent of point permutation. Following the constraint definitions in CAD software is not applicable. However, once the constraint is applied, the positional relations between primitives become fixed. Therefore, if the type of primitives and their relations are determined, the constraints on CAD shapes can be defined comprehensively.
The point-wise primitive type can be represented as the type of the primitive to which the point is attached, using one-hot encoding. The positional relation is defined as the combination of main axis direction and adjacency. Considering that the direction of primitives’ main axis is unique, as shown in Fig. 4, relations such as parallel and vertical could be derived from the direction of their main axes. However, some dimensional relations can not, such as the cylinders’ radius equivalence. Thus, additional data is required to supplement this information. Since the connected primitives are the basis of CAD shapes, the adjacency and main axis direction impose limitations on the distance and size of primitives. For example, in a cuboid, the adjacency and normal of the six faces ensure that the distance between the top and bottom faces equals the height of the front face, and the size of the back face is the same as the front one. Therefore, if the adjacency is combined with main axis directions, the positional relation can be determined comprehensively. Adjacency represents the connection between primitives and cannot be expressed as point-wise attribute directly. The edges are key indicators of adjacency, with each edge corresponding to two adjacent primitives, adjacency can therefore be inferred from edges. For point cloud, an edge can be represented by marking points near it, thus, point-wise adjacency is expressed as whether the point is near an edge, which can be represented by two-channel one-hot vector. In summary, each composition of constraints in parametric point cloud is illustrated in Fig. 5, and each point with constraint is expressed as (X, Y, Z, MAD, Adj, PT).
Figure 4: Main Axis Direction of primitives. Plane: normal vector. Cylinder, cone, sphere: rotation axis. Main Axis Direction requires unique process, details are provided in Supplementary.
We represent constraints by MAD, Adj, PT, is this approach sufficient to handle most constraints? For PT, commercial CAD systems typically classify primitives into plan, cylinder, cone, sphere, and free-form surface. Among these, plane, cylinder, and cone are primarily influenced by constraints, while spherical and free-form surfaces rely heavily on their geometric features. Therefore, considering plane, cylinder, and cone is sufficient for constraint representation. For sphere and free-form surface, the Stage 2 of CstNet also incorporate geometric features (feed in coordinates). MAD correspond to directional constraints (e.g. parallel and vertical), and MAD combined with PT represent dimensional constraints (e.g. distance and tangent). Through MAD and Adj, most common constraint types can be covered. Therefore, the combination of MAD, Adj, and PT is sufficient to represent the majority of constraints.
@@ 3.3 Constraint Acquisition
The Stage 1 of CstNet is designed to obtain constraints, as shown in the left of Fig. 7. When B-Rep data is available, the constraint computation from B-Rep (CstBRep) module is developed to obtain constraints. For the input B-Rep data, tessellation is performed by OCCT to generate a mesh model, followed by Poisson disk sampling [25] via pymeshlab to ensure a uniform point distribution and control the point count, if point cloud generation is not needed, this step could be skipped. For adjacency computation, we identify valid edges firstly. An edge is considered invalid if the two adjacent primitives are smoothly connected along the edge. Searching valid edges is accomplished by OCCT. Next, we compute the distance from each point to all valid edges. A point is marked as near an edge if the minimum distance is below a threshold (set to $0.08\times ShapeArea/4\pi$ in this paper). For the primitive type and main axis direction of a point, we first determine which primitive it attached, OCCT is then used to analyze the primitive, providing both the primitive type and main axis direction.
@@
In the absence of B-Rep data, such as working with mesh or voxel models, it can be converted to point cloud, after which the CstPnt module can be used to predict constraints. Since each point’s constraints are local information, the computation only requires data from its neighboring points, meaning that global information is unnecessary. The constraint computation of a point $\boldsymbol{p}$ could be divided into the following steps: Step 1: Identify neighbor points around $\boldsymbol{p}$. Step 2: From all neighbor points, identify which belongs to the same primitive as $\boldsymbol{p}$, store in $\boldsymbol{\mathrm{P_{valid}}}$. Step 3: Fit shapes such as cylinders or planes using $\boldsymbol{\mathrm{P_{valid}}}$, and determine the primitive type with the smallest fitting error. Step 4: Based on $\boldsymbol{\mathrm{P_{valid}}}$ and the primitive type, calculate the main axis direction and assess whether the point is near an edge. While the above process can be carried out using traditional methods, it is highly complex, which is why a deep learning model is employed to achieve the same.
@@
The design of CstPnt module is inspired by the above steps, with details in Fig. 7. For Step 1, the traditional approach is to use the k-Nearest Neighbors (KNN) algorithm, which identifies neighbors within a sphere. However, for constraint computation, it is more effective to search for neighbors along the shape’s surface. This approach captures more relevant points while reducing the influence of redundant points, as shown in Fig. 6. To this end, we proposed a neighbor searching method called SurfaceKNN, details in supplementary. For Step 2, the attention mechanism between points is employed to approximate this process, which applies different weights to neighbor points when updating the center point’s embedding, the relation between points is designed as embedding subtraction, as shown in Eq. 1. Step 3 and Step 4 involve computing MAD and PT, both steps can be accomplished by MLPs. We employed a U-Net-like structure for better data utilization.
@@
$\boldsymbol{f}^{\prime}_{i}=\sum_{f_{j}\in\mathcal{N}_{i}}\rho\Bigl{(}\rm{MLP}\bigl{(}\mathrm{Q}(\boldsymbol{f}_{\mathit{i}})-\mathrm{K}(\boldsymbol{f}_{\mathit{j}})\bigr{)}\Bigr{)}\odot\mathrm{V}(\boldsymbol{f}_{\mathit{i}}).$ (1)
@@
Where $\mathcal{N}_{i}$ is the feature collection of all neighbor points around $\boldsymbol{p}_{i}$, $\rho$ is the normalization function (softmax in this paper), and $\odot$ denotes element-wise multiplication. $\mathrm{Q}(\boldsymbol{f})$, $\mathrm{K}(\boldsymbol{f})$, and $\mathrm{V}(\boldsymbol{f})$ are defined in similar structure by concatenating $\boldsymbol{f}$ with the positional encoding $\Delta$, and then feeding it into MLPs, where $\Delta$ is defined as $\Delta=\boldsymbol{p}_{i}-\boldsymbol{p}_{j}$.
Figure 6: Comparison of KNN and SurfaceKNN. Blue point: center of neighbor search, red point: searched neighbors.
@@
The structure of CstPnt reveals that only local information is utilized, which aligns with the local nature of the constraint representation we designed. For no global information is incorporated and most CAD shapes consist of a limited set of primitives, such as plane, cylinder, and sphere, the CstPnt could generalize to unseen datasets if pre-trained on a large scale dataset. The training data is derived from unlabeled B-Rep data, as a result, our CstPnt enables the effective use of existing large scale unlabeled B-Rep datasets, making it possible to serve meaningful tasks.
Figure 7: Overview of CstNet. The top side illustrates the overall architecture of CstNet, which comprising two stages. Stage 1 is designed for constraint acquisition. When B-Rep data is available, the CstBRep module is utilized to extract constraints; otherwise, use the CstPnt module. Stage 2 performs constraint feature learning, facilitating a deeper understanding of CAD shapes. The bottom side presents the details of module design. In the CstPnt module, the final MLP consists of three parallel MLPs with the same input, where the outputs correspond to Main Axis Direction, Adjacency, and Primitive Type.
3.4 Constraint-Aware Feature Learning